# Testing Noise Assumptions of Learning Algorithms

### v1

@@ 2 Polynomial-Time Tester-Learners
Notation. We denote with $\mathbb{R},\mathbb{N},\mathbb{Z}$ the sets of real, natural and integer numbers correspondingly. For simplicity, we denote the $d$-dimensional standard Gaussian distribution as $\mathcal{N}_{d}$ or $\mathcal{N}$ if $d$ is clear by the context. For any set $S$, let $\operatorname{Unif}(S)$ denote the uniform distribution over $S$. We may also use the notation $\mathbf{x}\sim S$ in place of $\mathbf{x}\sim\operatorname{Unif}(S)$. For a set of points in $\mathbb{R}^{d}$, we denote with $\bar{S}$ the corresponding labeled set over $\mathbb{R}^{d}\times\{\pm 1\}$ where the corresponding labels will be clear by the context. For a vector $\mathbf{x}\in\mathbb{R}^{d}$, we denote with $x_{i}$ its $i$-th coordinate, i.e., $x_{i}=\mathbf{x}\cdot\mathbf{e}_{i}$.
@@ 2.1 Warm-up: Random Classification Noise Oracles
Relating the output error to optimum error. Let $\bar{S}$ be the input set of labeled examples and let $\mathbf{v}\in\mathbb{S}^{d-1}$ be the output of $\mathcal{A}$ on input $\bar{S}$. Note that, since $\bar{S}$ is not necessarily generated by $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},f^{*},\eta_{0}}$ (according to Definition 1.3), we do not have any a priori guarantees on $\mathbf{v}$. We may relate the output error $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})]$ to the optimum error $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})]$, by accounting for the set $\bar{S}_{g}$ of points $(\mathbf{x},y)\in\bar{S}$ that are labeled correctly by $\mathbf{v}$ but incorrectly by $\mathbf{v}^{*}$, as well as the set $\bar{S}_{b}$ of points in $\bar{S}$ that are labeled incorrectly by $\mathbf{v}$ but correctly by $\mathbf{v}^{*}$. Overall, we have the following
@@
Figure 1: The shaded region is $\{\mathbf{x}\in\mathbb{R}^{d}:\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})\}$. Left: red square points have label $+1$, blue round points have label $-1$. Right: green square points are in $\bar{S}_{g}$ and purple round points are in $\bar{S}_{b}$.
@@
The term ${|\bar{S}_{\mathsf{False}}|}/{|\bar{S}|}$ can be explicitly computed, since we have $\mathbf{v}$ and $\bar{S}$ and we can verify whether its value is at most $1/2-c$, as would be the case if the noise was RCN. Otherwise, we may safely reject. Now, we want to obtain certificates that the first term in Equation 2.4 can’t be too large and the second term can’t be too small.
The disagreement tester and how it is applied. If the example oracle was indeed in $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},\mathcal{H}_{\mathsf{hs}},\eta_{0}}$, then both probabilities in Equation 2.4 would equal to ${\measuredangle(\mathbf{v},\mathbf{v}^{*})}/{\pi}$. To see this, first observe that in the RCN case we would have $\mathbf{v}=\mathbf{v}^{*}$, $\mathbf{x}\sim\mathcal{N}_{d}$ and, moreover, even conditioned on the event $y\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})$, $\mathbf{x}$ remains Gaussian. We also use the following basic fact.
@@
It turns out that this property can be approximately certified by an efficient tester if we allow for a small, constant multiplicative slackness. The factor $1/2-c$ is crucial in order to be able to withstand this multiplicative angle amplification. Formally, we prove the existence of the following tester.
@@
Designing the disagreement tester. Our disagreement tester builds on ideas from prior work on testable agnostic learning by [GKSV24]. In particular, [GKSV24] show that when the angle between the input vector $\mathbf{v}$ and some unknown vector $\mathbf{v}^{\prime}$ is $\epsilon$, then one can give a testable bound of $O(\epsilon)$ on the quantity $\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})]$ by running some efficient tester (Proposition D.1 in [GKSV24]). To achieve this, the region $\{\mathbf{x}\in\mathbb{R}^{d}:\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})\}$ is covered by a disjoint union of simple regions whose masses can be testably upper bounded. In order to bound the mass of the simple regions, it is crucial to use the fact that $\mathbf{v}$ is known.
Here, our approach needs to be more careful, since we (1) require both upper and lower bounds on the quantity $\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})]$, (2) we do not have a specific target threshold for the angle $\measuredangle(\mathbf{v},\mathbf{v}^{\prime})$, but we need to provide testable bounds that involve $\measuredangle(\mathbf{v},\mathbf{v}^{\prime})$ as a free parameter and (3) we can only tolerate a small constant multiplicative error factor $(1\pm\mu)$. To obtain this improvement, we combine the approach of [GKSV24] with the notion of sandwiching polynomial approximators. Sandwiching polynomial approximators are also used to design testable learning algorithms (see [GKK23]), but we use them here in a more specialized way, by allowing the sandwiching function to be piecewise-polynomial.
@@
Moment matching and Chow matching in particular are known to have applications in testable learning (see, e.g., [GKK23, RV23, GKSV24, KSV24b, CKK+24]), but here we have to use this tool in a careful way. More specifically, we prove the following lemma (see Proposition B.6) based on a delicate argument that uses the sandwiching approximators of [DGJ+10, GOWZ10] (see Section B.3).
@@
Based on the above lemma (and a symmetric argument for the case $\mathbf{v}\cdot\mathbf{x}<0\leq\mathbf{v}^{\prime}\cdot\mathbf{x}$), the disagreement tester of Theorem 2.4 only needs to test constant-degree Chow parameters of the functions $\mathcal{I}_{i}(\mathbf{x})$, as described in Algorithm 1. Due to standard concentration arguments, if $S$ was i.i.d. from the Gaussian distribution, then the tests would pass.
@@
For any $\epsilon^{\prime},\delta\in(0,1)$ and $\eta_{0}=1/2-c$, where $c>0$ is any constant, there is an algorithm with time and sample complexity $\mathrm{poly}(d,1/\epsilon)\log(1/\delta)$ that has access to an example oracle $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},f^{*},\eta_{0}}$ for some unknown $f^{*}(\mathbf{x})=\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})$, $\mathbf{v}^{*}\in\mathbb{S}^{d-1}$ and outputs $\mathbf{v}\in\mathbb{S}^{d-1}$ such that $\measuredangle(\mathbf{v}^{*},\mathbf{v})\leq\epsilon^{\prime}$, with probability at least $1-\delta$.
@@ 2.2 Massart Noise Oracles
We now turn to the more challenging task of testable learning with respect to Gaussian Massart oracles and state out main theorem which shows that there is a fully polynomial-time algorithm even in this case. Recall that, informally, for some ground-truth halfspace $f$, the Massart oracle $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}$ outputs a Gaussian example $\mathbf{x}$, and with probability $\eta_{0}$ the adversary is given an option to make the accompanying label incorrect (i.e. $-f(\mathbf{x})$), and otherwise the label is correct (i.e. $f(\mathbf{x})$). The Massart noise model is known to be more challenging that the RCN model, because the label noise (in general) does not have symmetry properties that can be harnessed to make error terms coming from different regions cancel each other out.
@@
Fortunately, the first difference is not an issue, since appropriate results are known from prior work on classical learning under Massart noise and Gaussian marginal (see, e.g., [ABL17, DKTZ20a]).
Fact 2.8 ([DKTZ20a]).
@@
The spectral tester and how it is applied. In order to resolve the second complication and provide a certificate bounding the quantity $|\bar{S}_{b}|/|\bar{S}|$, we follow a different testing approach, based on ideas from tolerant testable learning [GSSV24], where the testers must accept whenever the input distribution is close to the target (and not necessarily equal). In particular, the spectral techniques of [GSSV24] are very robust to deletions. We provide the following tester which is guaranteed to accept subsets of Gaussian samples, since it is monotone under datapoint removal.
@@
Designing the spectral tester. Theorem 2.9 follows from a combination of ideas used to prove Theorem 2.4 and the spectral testing approach of [GSSV24]. In particular, instead of matching the Chow parameters $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[\mathcal{I}_{i}(\mathbf{x})\prod_{j\in[d]}\mathbf{x}^{\alpha_{i}}]$ of the quantities $\mathcal{I}_{i}$ as in Algorithm 1, we bound the maximum singular value of the Chow parameter matrices $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(\mathbf{x}^{\otimes k})(\mathbf{x}^{\otimes k})^{\top}\mathcal{I}_{i}(\mathbf{x})]$, where $\mathbf{x}^{\otimes k}$ denotes the vector of monomials of degree at most $k$. This can be done efficiently via the SVD algorithm and, crucially, satisfies the monotonicity under removal property of Theorem 2.9. Moreover, once the Chow parameter matrix is bounded, we have a bound for all quantities of the form $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}\mathcal{I}_{i}(\mathbf{x})]$, where $p$ is of degree at most $k$. Combining this observation with an analysis similar to the one for Lemma 2.5 (see Propositions C.5 and C.4) and a stronger version of the sandwiching polynomials of [DGJ+10, GOWZ10] by [KSV24b], we obtain that Algorithm 2 satisfies Theorem 2.9.
Overall algorithm. The overall algorithm receives an input set of labeled examples $\bar{S}$ and obtains a candidate $\mathbf{v}\in\mathbb{S}^{d-1}$ by running the algorithm of [DKTZ20a] with parameter $\epsilon^{\prime}=\epsilon^{3/2}/(C\sqrt{d})$ for some large enough constant $C$ (see 2.8). Then, it runs the disagreement tester of Theorem 2.4 with parameters $(S,\mathbf{v},\epsilon,\delta,\mu)$ (for some small constant $\mu$ depending on the noise rate $\eta_{0}=1/2-c$). Subsequently, the tester checks whether $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})]$ is at most $(1/2-c/2)|\bar{S}|$.
@@
$\bar{S}_{\mathsf{False}}^{\mathsf{far}}$ =$=\Bigr{\{}(\mathbf{x},y)\in\bar{S}_{\mathsf{False}}:|\measuredangle(\mathbf{v},\mathbf{x})-{\pi}/{2}|>{\epsilon^{3/2}}/{(d-1)^{1/2}}\Bigr{\}}\;\;\;\text{ and }\;\;\;\bar{S}_{\mathsf{False}}^{\mathsf{near}}=\bar{S}_{\mathsf{False}}\setminus\bar{S}_{\mathsf{False}}^{\mathsf{far}}$

### v2

@@ 2 Polynomial-Time Tester-Learners
Notation. We denote with $\mathbb{R},\mathbb{N},\mathbb{Z}$ the sets of real, natural and integer numbers correspondingly. For simplicity, we denote the $d$-dimensional standard Gaussian distribution as $\mathcal{N}_{d}$ or $\mathcal{N}$ if $d$ is clear by context. For any set $S$, let $\operatorname{Unif}(S)$ denote the uniform distribution over $S$. We may also use the notation $\mathbf{x}\sim S$ in place of $\mathbf{x}\sim\operatorname{Unif}(S)$. For a set of points in $\mathbb{R}^{d}$, we denote with $\bar{S}$ the corresponding labeled set over $\mathbb{R}^{d}\times\{\pm 1\}$ where the corresponding labels are those in the input of the algorithm unless otherwise specified. For a vector $\mathbf{x}\in\mathbb{R}^{d}$, we denote with $x_{i}$ its $i$-th coordinate, i.e., $x_{i}=\mathbf{x}\cdot\mathbf{e}_{i}$.
@@ 2.1 Warm-up: Random Classification Noise Oracles
Relating the output error to optimum error. Let $\bar{S}$ be the input set of labeled examples and let $\mathbf{v}\in\mathbb{S}^{d-1}$ be the output of $\mathcal{A}$ on input $\bar{S}$. Note that, since $\bar{S}$ is not necessarily generated by $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},f^{*},\eta_{0}}$, we do not have any a priori guarantees on $\mathbf{v}$. We may relate the output error $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})]$ to the optimum error $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})]$, by accounting for the set $\bar{S}_{g}$ of points $(\mathbf{x},y)\in\bar{S}$ that are labeled correctly by $\mathbf{v}$ but incorrectly by $\mathbf{v}^{*}$, as well as the set $\bar{S}_{b}$ of points in $\bar{S}$ that are labeled incorrectly by $\mathbf{v}$ but correctly by $\mathbf{v}^{*}$. Overall, we have the following
@@
Figure 1: The shaded region is $\{\mathbf{x}\in\mathbb{R}^{d}:\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})\}$. Left: red square points have label +1+1, blue round points have label 1-1. Right: green square points are in $\bar{S}_{g}$ and purple round points are in $\bar{S}_{b}$.
@@
The term ${|\bar{S}_{\mathsf{False}}|}/{|\bar{S}|}$ can be explicitly computed, since we have $\mathbf{v}$ and $\bar{S}$ and we can verify whether its value is at most 1/2c1/2-c, as would be the case if the noise was RCN. Otherwise, we may safely reject. Now, we want to obtain certificates that the first term in Equation 2.4 can’t be too large and the second term can’t be too small.
The disagreement tester and how it is applied. Our goal is to certify that both probabilities in Equation 2.4 are approximately equal to ${\measuredangle(\mathbf{v},\mathbf{v}^{*})}/{\pi}$, which is what we would expect if the example oracle were indeed in $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},\mathcal{H}_{\mathsf{hs}},\eta_{0}}$. This is because of the following fact, as well as the fact that even after conditioning on the event $y\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})$, $\mathbf{x}$ remains Gaussian.
@@
Recall, however, that we do not make any assumptions on the input examples. Therefore, we would like to certify the guarantee of 2.3. We show that this is possible by developing the following tester.
@@
Designing the disagreement tester. Our disagreement tester builds on ideas from prior work on testable agnostic learning by [gollakota2023efficient]. In particular, [gollakota2023efficient] show that when the angle between the input vector $\mathbf{v}$ and some unknown vector $\mathbf{v}^{\prime}$ is $\epsilon$, then one can give a testable bound of $O(\epsilon)$ on the quantity $\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})]$ by running some efficient tester (Proposition D.1 in [gollakota2023efficient]). To achieve this, the region $\{\mathbf{x}\in\mathbb{R}^{d}:\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})\}$ is covered by a disjoint union of simple regions whose masses can be testably upper bounded. In order to bound the mass of the simple regions, it is crucial to use the fact that $\mathbf{v}$ is known.
Here, our approach needs to be more careful, since we (1) require both upper and lower bounds on the quantity $\operatorname*{\mathbb{P}}_{\mathbf{x}\sim S}[\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})\neq\operatorname{\mathsf{sign}}(\mathbf{v}^{\prime}\cdot\mathbf{x})]$, (2) we do not have a specific target threshold for the angle $\measuredangle(\mathbf{v},\mathbf{v}^{\prime})$, but we need to provide testable bounds that involve $\measuredangle(\mathbf{v},\mathbf{v}^{\prime})$ as a free parameter and (3) we can only tolerate a small constant multiplicative error factor $(1\pm\mu)$. To obtain this improvement, we combine the approach of [gollakota2023efficient] with the notion of sandwiching polynomial approximators. Sandwiching polynomial approximators are also used to design testable learning algorithms (see [gollakota2022moment]), but we use them here in a more specialized way, by allowing the sandwiching function to be piecewise-polynomial.
@@
Moment matching and Chow matching in particular are known to have applications in testable learning (see, e.g., [gollakota2022moment, rubinfeld2022testing, gollakota2023efficient, klivans2023testable, chandrasekaran2024efficient]), but here we have to use this tool in a careful way. More specifically, we prove the following lemma (see Proposition B.6) based on a delicate argument that uses the sandwiching approximators of [diakonikolas2010bounded, gopalan2010fooling] (see Section B.3).
@@
Based on the above lemma (and a symmetric argument for the case $\mathbf{v}\cdot\mathbf{x}<0\leq\mathbf{v}^{\prime}\cdot\mathbf{x}$), the disagreement tester of Theorem 2.4 only needs to test quantities of the form $\operatorname*{\mathbb{E}}[\mathcal{I}_{i}\prod_{j}x_{j}^{\alpha_{j}}]$, which are known as constant-degree Chow parameters [chow-parameters, o2008chow] of the functions $\mathcal{I}_{i}(\mathbf{x})$, as described in Algorithm 1. Due to standard concentration arguments, if $S$ was i.i.d. from the Gaussian distribution, then the tests would pass.
@@
For any $\epsilon^{\prime},\delta\in(0,1)$ and $\eta_{0}=1/2-c$, where $c>0$ is any constant, there is an algorithm with time and sample complexity $\mathrm{poly}(d,1/\epsilon^{\prime})\log(1/\delta)$ that has access to an example oracle $\mathsf{EX}^{\mathsf{RCN}}_{\mathcal{N},f^{*},\eta_{0}}$ for some unknown $f^{*}(\mathbf{x})=\operatorname{\mathsf{sign}}(\mathbf{v}^{*}\cdot\mathbf{x})$, $\mathbf{v}^{*}\in\mathbb{S}^{d-1}$ and outputs $\mathbf{v}\in\mathbb{S}^{d-1}$ such that $\measuredangle(\mathbf{v}^{*},\mathbf{v})\leq\epsilon^{\prime}$, with probability at least $1-\delta$.
@@ 2.2 Massart Noise Oracles
We now turn to the more challenging task of testable learning with respect to Gaussian Massart oracles and state our main theorem which shows that there is a fully polynomial-time algorithm even in this case. Observe that, informally, for some ground-truth halfspace $f$, the Massart oracle $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}$ can be equivalently viewed as follows: The oracle outputs a Gaussian example $\mathbf{x}$, and with probability $\eta_{0}$ the adversary is given an option to make the accompanying label incorrect (i.e. $f(\mathbf{x})$), and otherwise the label is correct (i.e. $f(\mathbf{x})$). The Massart noise model is known to be more challenging that the RCN model, because the label noise (in general) does not have symmetry properties that can be harnessed to make error terms coming from different regions cancel each other out.
@@
Fortunately, the first difference is not an issue, since appropriate results are known from prior work on classical learning under Massart noise and Gaussian marginal (see, e.g., [awasthi2017power, diakonikolas2020learning]).
Fact 2.8 ([diakonikolas2020learning]).
@@
The spectral tester and how it is applied. In order to resolve the second complication and provide a certificate bounding the quantity $|\bar{S}_{b}|/|\bar{S}|$, we follow a different testing approach, based on ideas from tolerant testable learning [goel2024tolerant], where the testers must accept whenever the input distribution is close to the target (and not necessarily equal). We provide the following tester which is guaranteed to accept subsets of Gaussian samples, since it is monotone under datapoint removal.
@@
Designing the spectral tester. Theorem 2.9 follows from a combination of ideas used to prove Theorem 2.4 and the spectral testing approach of [goel2024tolerant]. In particular, instead of matching the Chow parameters $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[\mathcal{I}_{i}(\mathbf{x})\prod_{j\in[d]}\mathbf{x}^{\alpha_{i}}]$ of the quantities $\mathcal{I}_{i}$ as in Algorithm 1, we bound the maximum singular value of the Chow parameter matrices $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(\mathbf{x}^{\otimes k})(\mathbf{x}^{\otimes k})^{\top}\mathcal{I}_{i}(\mathbf{x})]$, where $\mathbf{x}^{\otimes k}$ denotes the vector of monomials of degree at most $k$. This can be done efficiently via the SVD algorithm and, crucially, satisfies the monotonicity under removal property of Theorem 2.9. Moreover, once the Chow parameter matrix is bounded, we have a bound for all quantities of the form $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}\mathcal{I}_{i}(\mathbf{x})]$, where $p$ is of degree at most $k$. Combining this observation with an analysis similar to the one for Lemma 2.5 (see Propositions C.5 and C.4) and a stronger version of the sandwiching polynomials of [diakonikolas2010bounded, gopalan2010fooling] by [klivans2023testable], we obtain that Algorithm 2 satisfies Theorem 2.9.
Overall algorithm. The overall algorithm receives an input set of labeled examples $\bar{S}$ and obtains a candidate $\mathbf{v}\in\mathbb{S}^{d-1}$ by running the algorithm of [diakonikolas2020learning] with parameter $\epsilon^{\prime}=\epsilon^{3/2}/(C\sqrt{d})$ for some large enough constant $C$ (see 2.8). Then, it runs the disagreement tester of Theorem 2.4 with parameters $(S,\mathbf{v},\epsilon,\delta,\mu)$ (for some small constant $\mu$ depending on the noise rate $\eta_{0}=1/2-c$). Subsequently, the tester checks whether $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq\operatorname{\mathsf{sign}}(\mathbf{v}\cdot\mathbf{x})]$ is at most 1/2c/21/2-c/2.
@@
$\bar{S}_{\mathsf{False}}^{\mathsf{far}}$ $=\Bigr\{(\mathbf{x},y)\in\bar{S}_{\mathsf{False}}:|\measuredangle(\mathbf{v},\mathbf{x})-{\pi}/{2}|>{\epsilon^{3/2}}/{(d-1)^{1/2}}\Bigr\}\;\;\;\text{ and }\;\;\;\bar{S}_{\mathsf{False}}^{\mathsf{near}}=\bar{S}_{\mathsf{False}}\setminus\bar{S}_{\mathsf{False}}^{\mathsf{far}}$